40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 13 00:56:07 2016, gautam, Update, endtable upgrade, X endtable repopulation 
    Reply  Wed Apr 13 11:00:28 2016, Steve, Update, endtable upgrade, X endtable repopulation 
       Reply  Wed Apr 13 18:25:07 2016, gautam, Update, endtable upgrade, Lightwave health check 
          Reply  Fri Apr 15 18:38:12 2016, gautam, Update, endtable upgrade, Lightwave health check - NO IMPROVEMENT 
             Reply  Tue Apr 19 18:37:29 2016, gautam, Update, endtable upgrade, Laser swap + optical layout IR_modematch_19April2016_2.pdf
                Reply  Thu Apr 21 14:25:52 2016, gautam, Update, endtable upgrade, Green light recovered IMG_6567.JPGIMG_6568.JPGCVI_reflectivity.jpeg
                Reply  Tue Apr 26 15:22:35 2016, Steve, Update, endtable upgrade, Cleaning ETMX vacuum dirty window ETMX-Tvp.jpgETMX-TvpDetail.jpg
                   Reply  Tue Apr 26 23:19:42 2016, gautam, Update, endtable upgrade, Green aligned to arm - high order mode flashes seen IMG_2219.JPGIMG_2220.JPG
                      Reply  Thu Apr 28 00:41:08 2016, gautam, Update, endtable upgrade, more progress - Transmon PD installed 
                         Reply  Fri Apr 29 00:55:46 2016, gautam, Update, endtable upgrade, green PDH locked to Xarm 
                            Reply  Mon May 2 19:14:18 2016, gautam, Update, endtable upgrade, Optical layout almost complete ETMX_wedge.pdf
                               Reply  Thu May 5 03:05:37 2016, gautam, Update, endtable upgrade, ALS status update ALS_noiseSpec_5May2016_2.pdfCoherence_5May2016.pdfimage.jpeg
                                  Reply  Thu May 5 04:05:03 2016, ericq, Update, LSC, Aux X PDH checks nom_noise.pngnom_olg.png
                                     Reply  Thu May 5 14:03:52 2016, ericq, Update, LSC, Further Aux X PDH tweaks 2016-05-05_newals.pdf
                                  Reply  Thu May 5 14:05:01 2016, rana, Update, endtable upgrade, ALS status update 
                                  Reply  Thu May 5 21:28:44 2016, gautam, Update, endtable upgrade, Innolight PZT capacitance 
                               Reply  Thu May 19 16:29:20 2016, Steve, Update, endtable upgrade, Optical layout almost complete ETMX_4x3_closed.jpgsealedETMXenclosure.jpg
                                  Reply  Fri Sep 30 19:53:07 2016, gautam, Update, endtable upgrade, X end IR pickoff fiber coupled 
                                     Reply  Wed Oct 5 16:28:10 2016, gautam, Update, endtable upgrade, EX laser power monitor PD installed 
                   Reply  Wed Apr 27 15:04:47 2016, Steve, Update, endtable upgrade, Cleaning ETMX vacuum dirty window 1cETMX-Tcmp.jpg
                      Reply  Fri Apr 29 16:05:23 2016, gautam, Update, endtable upgrade, Cleaning ETMX vacuum dirty window IMG_0755.JPG
Message ID: 12095     Entry time: Thu Apr 28 00:41:08 2016     In reply to: 12090     Reply to this: 12099
Author: gautam 
Type: Update 
Category: endtable upgrade 
Subject: more progress - Transmon PD installed 

The IR Transmon system is almost completely laid out, only the QPD remains to be installed. Some notes:

  1. The "problem" with excessive green power reflected from the harmonic separator has been resolved. It is just very sensitive to the angle of incidence. In the present configuration, there is ~10uW of green power reflected from either side, which shouldn't be too worrisome. But this light needs to be dumped. Given the tiny amount, I think a black glass + sticky tape solution is best suited, given the space constraints. This does not reach the Transmon PDs because there is a filter in the path that is transmissive to IR only. 
  2. I aligned the transmitted beam onto the Thorlabs PD, and reconnected the signal BNC cable (the existing cable wasn't long enough so I had to use a barrel connector and a short extension cable). I then reverted the LSC trigger for the X arm back to TRX DC and also recompiled c1ass to revert to TRX for the dither alignment. At the moment, both arms are stably locked, although the X arm transmission is saturated at ~0.7 after running the dither alignment. I'm not sure if this is just a normalization issue given the new beam path or if there is something else going on. Further investigations tomorrow.
  3. It remains to dump some of the unwanted green light from the addition of the harmonic separator...
  4. We may want to redesign some (or all) of the Transmon path - the lens currently in use seems to have been chosen arbitrarily. Moreover, it is quite stubbornly dirty, there are some markings which persist after repeated first contact cleaning...

I feel like once the above are resolved, the next step would be to PDH lock the green to the arm and see what sort of transmission we get on the PSL table. It may be the polarization or just alignment, but for some reason, the transmitted green light from the X arm is showing up at GTRY now (up to 0.5, which is the level we are used to when the Y arm has green locked!). So a rough plan of action:

  1. Install transmon QPD
  2. PDH lock green to X arm
  3. Fix the window situation - as Steve mentioned in an earlier elog, the F.C. cleaning seems to have worked well, but a little remains stuck on the window (though away from where any laser radiation is incident). This is resolved easily enough if we apply one more layer of F.C., but the bottle-neck right now is we are out of PEEK which is what we use to remove the F.C. once dried. Steve thinks a fresh stock should be here in the next couple of days...
  4. Once 3 is resolved, we can go ahead and install the Oplev.
  5. Which leaves the lst subsystem, coupling to the fiber and a power monitor for the NPRO. I have resolved to do both these using the 1% transmitted beam after the beamsplitter immediately after the NPRO rather than pick off at the harmonic separator after the doubling oven. I need to do the mode-matching calculation for coupling into the fiber and also adjust the collimating lens...
  6. Clean-up: make sure cables are tied down, strain-relieved and hooked up to whatever they are supposed to be hooked up to...
ELOG V3.1.3-